# Validity of relativistic hydrodynamics beyond local equilibrium

### v1

@@ A.0.1 The G-Function
Appendix B 0+10+1D Calculations
B.1 Exact solutions to the moment equation
@@
We define the Operator
@@
which gives us the integrating factor,
@@
We now define the operator that we call the generator of Free streaming
@@
where $(\cdot)^{(k)}$ is the rising poccamer symbol.
@@
With the generator of hydrodynamics,
$\rho^{G}_{n,l}(\tau)$ $=\int_{\tau_{0}}^{\tau}d\tau^{\prime}\,\sum_{k=0}^{\infty}\frac{\left(l-\frac{n}{2}\right)^{(k)}}{k!}\frac{e^{-\xi^{\prime}}}{\tau_{R}}\left(\frac{\tau^{\prime}}{\tau}\right)^{2l+1}\left[\left(\frac{\tau^{\prime}}{\tau}\right)^{2}-1\right]^{k}\rho_{n,l+k}^{\text{eq}}(\tau^{\prime})$ (123)
We can now use the expression for $\rho_{n}^{\text{eq}}$,
$\rho^{G}_{n,l}(\tau)$ $=\int_{\tau_{0}}^{\tau}d\tau^{\prime}\,\sum_{k=0}^{\infty}\frac{\left(l-\frac{n}{2}\right)^{(k)}}{k!}\frac{e^{-\xi^{\prime}}}{\tau_{R}}\left(\frac{\tau^{\prime}}{\tau}\right)^{2l+1}\left[\left(\frac{\tau^{\prime}}{\tau}\right)^{2}-1\right]^{k}\frac{T^{n+3}}{2\pi^{2}}\frac{G_{n+3,l+k}(z)}{(2(l+k)+1)}e^{\mu/T}$ (124)
@@
$=\frac{\Gamma(n+3)}{(2\pi^{2})(2l+1)}\int_{\tau_{0}}^{\tau}d\tau^{\prime}\frac{e^{-\xi^{\prime}}}{\tau_{R}}\left(\frac{\tau^{\prime}}{\tau}\right)^{2l+1}F(l+\frac{1}{2},l-\frac{n}{2};l+\frac{3}{2};\left(\frac{\tau^{\prime}}{\tau}\right)^{2}-1)\,T^{n+3}$ (126)
B.2 0+10+1D Gradient expansion
@@
B.3 Hydrodynamics
Hydrodynamics in 0+10+1D requires us to find the dynamical equations for the energy density $\epsilon$ and the two anisotropies, bulk pressure $\Pi$ and shear pressure $\pi$. From energy conservation we get the dynamical equation for energy density ($\epsilon=\rho_{1,1}$),
$\mathbf{\hat{D}}\rho_{1,0}=\mathbf{\hat{D}}\text{n}=0$ (134)
$\mathbf{\hat{D}}\rho_{1,1}=\mathbf{\hat{D}}\epsilon=0$ (135)
@@ B.3 Hydrodynamics
$\partial_{\tau}\epsilon$ $=-\frac{\epsilon+P_{L}}{\tau}$ (137)
@@
$\mathbf{\hat{D}}\pi_{n,l}^{G}=-\frac{\pi_{n,l}^{G}}{\tau_{R}}-\mathbf{\hat{D}}\rho^{eq}_{n,l}$ (138)
@@
$\Pi$ $=\frac{1}{3}(\epsilon-3P)-\frac{1}{3}\langle m^{2}\rangle\,.$ (140)
@@
$\Pi$ $=\frac{1}{3}(\epsilon-3P)-\frac{1}{3}m^{2}(\rho_{-1,0}^{eq}+\pi_{-1,0})$
@@
We also need to compute the equilibrium gradient,
$\mathbf{\hat{D}}\rho^{eq}_{n,l}(\tau)=\partial_{\tau}\rho^{eq}_{n,l}(\tau)+\frac{(2l+1)}{\tau}\rho^{eq}_{n,l}(\tau)-\frac{(2l-n)}{\tau}\rho^{eq}_{1,2}(\tau)\,.$ (144)
@@
$\partial_{\tau}\rho^{eq}_{n,l}(\tau)$ $=\partialderivative{\rho^{eq}_{n,l}}{\epsilon}\partialderivative{\epsilon}{\tau}+\partialderivative{\rho^{eq}_{n,l}}{n}\partialderivative{n}{\tau}$ (145)
$=-\partialderivative{\rho^{eq}_{n,l}}{\epsilon}~\frac{(\epsilon+P+\Pi-\pi)}{\tau}-\partialderivative{\rho^{eq}_{n,l}}{\text{n}}~\frac{\text{n}}{\tau}\,.$ (146)
Define
$c^{2}_{n,l}$ $=\partialderivative{\rho^{eq}_{n,l}}{\epsilon}\,$, (147)
$\kappa_{n,l}$ $=n\partialderivative{\rho^{eq}_{n,l}}{\text{n}}\,.$ (148)
@@
$\partial_{\tau}\rho^{eq}_{n,l}(\tau)=-c_{n,l}^{2}\frac{(\epsilon+P+\Pi-\pi)}{\tau}-\frac{\kappa_{n,l}}{\tau}$ (149)
For a system with, we then obtain,
$\mathbf{\hat{D}}P$ $=-c_{s}^{2}\frac{(\epsilon+P+\Pi-\pi)}{\tau}-\frac{\kappa}{\tau}+\frac{3P}{\tau}-\frac{\rho^{eq}_{1,2}(\tau)}{\tau}$ (150)
Evolution of $\pi_{1,1}$
@@
$\mathbf{\hat{D}}\pi_{1,1}$ $=-\frac{\pi_{1,1}}{\tau_{R}}+c_{s}^{2}\frac{(\epsilon+P+\pi_{1,1})}{\tau}+\frac{\kappa}{\tau}-\frac{3P}{\tau}+\frac{\rho^{eq}_{1,2}(\tau)}{\tau}$ (151)
$=-\frac{\pi_{1,1}}{\tau_{R}}+c_{s}^{2}\frac{\pi_{1,1}}{\tau}+c_{s}^{2}\frac{(\epsilon+P)}{\tau}+\frac{\kappa}{\tau}-3\frac{P}{\tau}+\frac{\rho^{eq}_{1,2}(\tau)}{\tau}\,.$ (152)
We note that,

### v2

@@ A.0.1 The G-Function
We note the relation,
$z^{2}G_{n,l}(z)$ $=\int_{z}^{\infty}dyz^{2}y^{n-2l-2}(y^{2}-z^{2})^{l+1/2}e^{-y}$ (116)
$=-\int_{z}^{\infty}dy\left[(y^{2}-z^{2})-y^{2}\right]~y^{n-2l-2}(y^{2}-z^{2})^{l+1/2}e^{-y}$ (117)
$=-\int_{z}^{\infty}dyy^{n-2l-2}(y^{2}-z^{2})^{l+1+1/2}e^{-y}+\int_{z}^{\infty}dyy^{n+2-2l-2}(y^{2}-z^{2})^{l+1/2}e^{-y}$ (118)
$=G_{n+2,l}(z)-G_{n+2,l+1}(z)$ (119)
@@
Appendix B 0+10+1D Solution
@@
We define the generator of free streaming,
$\mathbf{\hat{F}}=\frac{2l+1}{\tau}\mathbf{\hat{I}}-\frac{2l-n}{\tau}\mathbf{\hat{S}}\,$. (123)
@@
To find the propagator of the solutions, we define,
@@
which gives us the integrating(propagator) factor,
@@
We now define the operator, which is the integral of the generator of Free streaming
@@
where $(\cdot)^{(k)}$ is the rising Pochhammer symbol.
@@
We define the kernel,
$K_{n,l,k}(\tau,\tau^{\prime})=\frac{\left(l-\frac{n}{2}\right)^{(k)}}{k!}\left(\frac{\tau^{\prime}}{\tau}\right)^{2l+1}\left[1-\left(\frac{\tau^{\prime}}{\tau}\right)^{2}\right]^{k}\,$, (139)
where $(\cdot)^{(}k)$ is the rising poccammer symbol. Note that
$K_{n,l,0}(\tau,\tau^{\prime})=\left(\frac{\tau^{\prime}}{\tau}\right)^{2l+1}\,$, (140)
and
$K_{n,l,k}(\tau,\tau)$ $=1$ $k=0\,$, (141)
$K_{n,l,k}(\tau,\tau)$ $=0$ $k\geq 1\,.$ (142)
@@
The generator of hydrodynamics is,
$\rho^{G}_{n,l}(\tau)$ $=\int_{\tau_{0}}^{\tau}d\tau^{\prime}\frac{e^{-\xi^{\prime}}}{\tau_{R}}\,\sum_{k=0}^{\infty}K_{n,l,k}(\tau,\tau^{\prime})\rho_{n,l+k}^{\text{eq}}(\tau^{\prime})$ (144)
We can now use the expression for $\rho_{n,l}^{\text{eq}}$,
$\rho^{G}_{n,l}(\tau)$ $=\int_{\tau_{0}}^{\tau}d\tau^{\prime}\,\frac{e^{-\xi^{\prime}}}{\tau_{R}}\,\sum_{k=0}^{\infty}K_{n,l,k}(\tau,\tau^{\prime})\frac{T^{n+3}}{2\pi^{2}}\frac{G_{n+3,l+k}(z)}{(2(l+k)+1)}e^{\mu/T}$ (145)
@@
$=\frac{\Gamma(n+3)}{(2\pi^{2})(2l+1)}\int_{\tau_{0}}^{\tau}d\tau^{\prime}\frac{e^{-\xi^{\prime}}}{\tau_{R}}\left(\frac{\tau^{\prime}}{\tau}\right)^{2l+1}F(l+\frac{1}{2},l-\frac{n}{2};l+\frac{3}{2};\left(\frac{\tau^{\prime}}{\tau}\right)^{2}-1)\,T^{n+3}\,$, (147)
where $F(a,b;c;x)$ is the hypergeometric function.
B.1 0+10+1D Gradient expansion
@@
Appendix C Hydrodynamic Equations
Hydrodynamics in 0+10+1D requires us to find the dynamical equations for the energy density $\varepsilon$ and the two anisotropies, bulk pressure $\Pi$ and shear pressure $\pi$. From energy conservation we get the dynamical equation for energy density ($\varepsilon=\rho_{1,0}$),
$\mathbf{\hat{D}}\rho_{0,0}=\mathbf{\hat{D}}\text{n}=0$ (155)
$\mathbf{\hat{D}}\rho_{1,0}=\mathbf{\hat{D}}\varepsilon=0$ (156)
@@ Appendix C Hydrodynamic Equations
$\partial_{\tau}\varepsilon$ $=-\frac{\varepsilon+P_{L}}{\tau}$ (158)
@@
$\mathbf{\hat{D}}\pi_{n,l}=-\frac{\pi_{n,l}}{\tau_{R}}-\mathbf{\hat{D}}\rho^{\rm eq}_{n,l}$ (159)
@@
$\Pi$ $=\frac{1}{3}(\varepsilon-3P)-\frac{1}{3}\langle m^{2}\rangle\,.$ (161)
@@
$\Pi$ $=\frac{1}{3}(\varepsilon-3P)-\frac{1}{3}m^{2}(\rho_{-1,0}^{\rm eq}+\pi_{-1,0})$
@@
In the conformal limit
$\pi=-\pi_{1,1}$ (165)
C.1 Speed of sound and bulk modulus
Define
$c^{2}_{n,l}$ $=\partialderivative{\rho^{\rm eq}_{n,l}}{\varepsilon}\,$, (166)
$\kappa_{n,l}$ $=n\partialderivative{\rho^{\rm eq}_{n,l}}{\text{n}}\,.$ (167)
with the speed of sound $c_{s}^{2}=c^{2}_{1,1}$ and bulk modulus $\kappa=\kappa_{1,1}$ as $\rho_{1,1}^{\rm eq}=P$, the equilibrium pressure.
We now find explicit expressions for the variables defined in Eq(166).
$\frac{\partial\rho_{n,l}^{eq}}{\partial T}$ $=\frac{\partial}{\partial T}\left[e^{\alpha}\,\frac{T^{n+3}}{(2l+1)2\pi^{2}}\,G_{n+3,l}(z)\right]$ (168)
$=e^{\alpha}\,\frac{(n+3)T^{2}}{(2l+1)2\pi^{2}}\,G_{n+3,l}\;+\;e^{\alpha}\,\frac{T^{3}}{(2l+1)2\pi^{2}}\,\frac{\partial G_{n+3,l}}{\partial z}\,\frac{\partial z}{\partial T}$ (169)
Using
$\frac{\partial z}{\partial T}=-\frac{z}{T},\qquad\frac{\partial G_{n+3,l}}{\partial z}=\frac{1}{z}((n+3)G_{n+3,l}-G_{n+4,l}),$
we get
$\frac{\partial G_{n+3,l}}{\partial T}=-\frac{1}{T}((n+3)G_{n+3,l}-G_{n+4,l})$ (170)
$\frac{\partial\rho_{n,l}^{eq}}{\partial T}$ $=e^{\alpha}\,\frac{(n+3)T^{n+2}}{(2l+1)2\pi^{2}}\,G_{n+3,l}-e^{\alpha}\,\frac{T^{n+3}}{(2l+1)2\pi^{2}}\,\frac{1}{T}\,((n+3)G_{n+3,l}-G_{n+4,l})$ (171)
$=e^{\alpha}\,\frac{T^{n+2}}{(2l+1)2\pi^{2}}G_{n+4,l}$ (172)
$\frac{\partial\rho_{n,l}^{eq}}{\partial\alpha}$ $=\frac{\partial}{\partial\alpha}\left[e^{\alpha}\,\frac{T^{n+3}}{(2l+1)2\pi^{2}}\,G_{n+3,l}\right]$ (173)
$=e^{\alpha}\,\frac{T^{n+3}}{(2l+1)2\pi^{2}}\,G_{n+3,l}\;=\;\rho_{n,l}^{eq}$ (174)
$\frac{\partial\varepsilon}{\partial T}=e^{\alpha}\,\frac{T^{3}}{2\pi^{2}}G_{5,0}$ $\frac{\partial\varepsilon}{\partial\alpha}=e^{\alpha}\,\frac{T^{4}}{2\pi^{2}}G_{4,0}$ (175)
$\frac{\partial\rm n}{\partial T}=e^{\alpha}\,\frac{T^{2}}{2\pi^{2}}G_{4,0}$ $\frac{\partial\rm n}{\partial\alpha}=e^{\alpha}\,\frac{T^{3}}{2\pi^{2}}G_{3,0}$ (176)
We then have the Jacobian relation,
$\begin{bmatrix}{\rm d}\varepsilon\\ {\rm d}\rm n\end{bmatrix}=e^{\alpha}\,\frac{T^{2}}{2\pi^{2}}\begin{bmatrix}TG_{5,0}&&T^{2}G_{4,0}\\ G_{4,0}&&TG_{3,0}\end{bmatrix}\begin{bmatrix}{\rm d}T\\ {\rm d}\alpha\end{bmatrix}$ (177)
The inverse Jacobian relation is
$\begin{bmatrix}{\rm d}T\\ {\rm d}\alpha\end{bmatrix}=e^{-\alpha}\,\frac{2\pi^{2}}{T^{4}}\frac{1}{G_{5,0}G_{3,0}-G_{4,0}^{2}}\begin{bmatrix}TG_{3,0}&&-T^{2}G_{4,0}\\ -G_{4,0}&&TG_{5,0}\end{bmatrix}\begin{bmatrix}{\rm d}\varepsilon\\ {\rm d}\rm n\end{bmatrix}$ (178)
From this we get the partial derivative relations
$\frac{\partial T}{\partial\varepsilon}=~~~e^{-\alpha}\,\frac{2\pi^{2}}{T^{3}}\frac{G_{3,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$ $\frac{\partial T}{\partial\rm n}=-e^{-\alpha}\,\frac{2\pi^{2}}{T^{2}}\frac{G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$ (179)
$\frac{\partial\alpha}{\partial\varepsilon}=-e^{-\alpha}\,\frac{2\pi^{2}}{T^{4}}\frac{G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$ $\frac{\partial\alpha}{\partial\rm n}=~~~~e^{-\alpha}\,\frac{2\pi^{2}}{T^{3}}\frac{G_{5,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$ (180)
Now we can find
$\frac{\partial\rho_{n,l}^{eq}}{\partial\varepsilon}$ $=\frac{\partial\rho_{n,l}^{eq}}{\partial T}\frac{\partial T}{\partial\varepsilon}+\frac{\partial\rho_{n,l}^{eq}}{\partial\alpha}\frac{\partial\alpha}{\partial\varepsilon}$ (181)
$=\left[e^{\alpha}\,\frac{T^{n+2}}{(2l+1)2\pi^{2}}G_{n+4,l}\right]\left(e^{-\alpha}\,\frac{2\pi^{2}}{T^{3}}\frac{G_{3,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}\right)$ (182)
$~~~~~~~~~~~~+\left[e^{\alpha}\,\frac{T^{n+3}}{(2l+1)2\pi^{2}}\,G_{n+3,l}\right]\left(-e^{-\alpha}\,\frac{2\pi^{2}}{T^{4}}\frac{G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}\right)$ (183)
$=\frac{T^{n-1}}{(2l+1)}\frac{G_{n+4,l}G_{3,0}-G_{n+3,l}G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$ (184)
$\frac{\partial\rho_{n,l}^{eq}}{\partial\rm n}$ $=\frac{\partial\rho_{n,l}^{eq}}{\partial T}\frac{\partial T}{\partial\rm n}+\frac{\partial\rho_{n,l}^{eq}}{\partial\alpha}\frac{\partial\alpha}{\partial\rm n}$ (185)
$=\left[e^{\alpha}\,\frac{T^{n+2}}{(2l+1)2\pi^{2}}G_{n+4,l}\right]~\left(-e^{-\alpha}\,\frac{2\pi^{2}}{T^{2}}\frac{G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}\right)$ (186)
$~~~~~~~~~~~~+\left[e^{\alpha}\,\frac{T^{n+3}}{(2l+1)2\pi^{2}}\,G_{n+3,l}\right]~\left(e^{-\alpha}\,\frac{2\pi^{2}}{T^{3}}\frac{G_{5,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}\right)$ (187)
$=\frac{T^{n}}{(2l+1)}\frac{G_{5,0}G_{n+3,l}-G_{4,0}G_{n+4,l}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$ (188)
so we find
$c_{n,l}^{2}$ $=\frac{T^{n-1}}{(2l+1)}\frac{G_{n+4,l}G_{3,0}-G_{n+3,l}G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}\,$, (189)
$\kappa_{n,l}$ $={\rm n}\frac{T^{n}}{(2l+1)}\frac{G_{5,0}G_{n+3,l}-G_{4,0}G_{n+4,l}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}\,$,. (190)
As a check note that $\frac{\partial\rho_{1,0}^{eq}}{\partial\varepsilon}=c_{1,0}^{2}=1$ and $\frac{\partial\rho_{0,0}^{eq}}{\partial\rm n}=\kappa_{0,0}=1$.
Finally, we get the speed of sound
$c_{s}^{2}(z)=c_{1,1}^{2}(z)=\frac{1}{3}\left[\frac{G_{5,1}(z)G_{3,0}(z)-G_{4,1}(z)G_{4,0}(z)}{G_{5,0}(z)G_{3,0}(z)-G_{4,0}^{2}(z)}\right]\,.$ (191)
In the conformal limit, we have
$c_{s}^{2}(0)=\frac{1}{3}\left[\frac{\Gamma(5)\Gamma(3)-\Gamma(4)\Gamma(4)}{\Gamma(5)\Gamma(3)-\Gamma(4)\Gamma(4)}\right]=\frac{1}{3}\,$,.
The bulk modulus,
$\kappa(z)=\kappa_{1,1}={\rm n}\frac{T}{(2l+1)}\frac{G_{5,0}(z)G_{4,1}(z)-G_{4,0}(z)G_{5,1}(z)}{G_{5,0}(z)G_{3,0}(z)-G_{4,0}^{2}(z)}\,$, (192)
which in the conformal limit
$\kappa(0)={\rm n}\frac{T}{(2l+1)}\frac{\Gamma(5)\Gamma(4)-\Gamma(4)\Gamma(5)}{\Gamma(5)\Gamma(3)-\Gamma(4)^{2}}=0\,$,
vanishes.
C.2 Evolution equations
We start with computing the equilibrium gradient,
$\mathbf{\hat{D}}\rho^{\rm eq}_{n,l}(\tau)=\partial_{\tau}\rho^{\rm eq}_{n,l}(\tau)+\frac{(2l+1)}{\tau}\rho^{\rm eq}_{n,l}(\tau)-\frac{(2l-n)}{\tau}\rho^{\rm eq}_{n,l+1}(\tau)\,.$ (193)
@@ C.2 Evolution equations
$\partial_{\tau}\rho^{\rm eq}_{n,l}(\tau)$ $=\partialderivative{\rho^{\rm eq}_{n,l}}{\varepsilon}\partialderivative{\varepsilon}{\tau}+\partialderivative{\rho^{\rm eq}_{n,l}}{n}\partialderivative{n}{\tau}$ (194)
$=-\partialderivative{\rho^{\rm eq}_{n,l}}{\varepsilon}~\frac{(\varepsilon+P+\Pi-\pi)}{\tau}-\partialderivative{\rho^{\rm eq}_{n,l}}{\text{n}}~\frac{\text{n}}{\tau}\,.$ (195)
@@
$\partial_{\tau}\rho^{\rm eq}_{n,l}(\tau)=-c_{n,l}^{2}\frac{(\varepsilon+P+\Pi-\pi)}{\tau}-\frac{\kappa_{n,l}}{\tau}$ (196)
$\mathbf{\hat{D}}\rho^{\rm eq}_{n,l}(\tau)$ $=-c_{n,l}^{2}\frac{(\varepsilon+P+\Pi-\pi)}{\tau}-\frac{\kappa_{n,l}}{\tau}+\frac{(2l+1)}{\tau}\rho^{\rm eq}_{n,l}(\tau)-\frac{(2l-n)}{\tau}\rho^{\rm eq}_{n,l+1}(\tau)$ (197)
$=-c_{n,l}^{2}\frac{\pi_{1,1}}{\tau}-\left[c_{n,l}^{2}(\varepsilon+P)+\kappa_{n,l}-\frac{3}{G_{n+3,1}(z)}\left(G_{n+3,l}(z)-\frac{(2l-n)}{(2l+3)}G_{n+3,l+1}(z)\right)\rho_{n,1}^{\rm eq}\right]\frac{1}{\tau}\,.$ (198)
For the equilibrium pressure $\rho_{1,1}^{\rm eq}$then obtain,
$\mathbf{\hat{D}}P$ $=-c_{s}^{2}\frac{(\varepsilon+P+\Pi-\pi)}{\tau}-\frac{\kappa}{\tau}+\frac{3P}{\tau}-\frac{\rho^{\rm eq}_{1,2}(\tau)}{\tau}$ (199)
$=-c_{s}^{2}\frac{\pi_{1,1}}{\tau}-\left(c_{s}^{2}(\varepsilon+P)+\kappa-3\left[1-\frac{1}{5}\frac{G_{4,2}(z)}{G_{4,1}(z)}\right]P\right)\frac{1}{\tau}$ (200)
The from (159) we get the evolution equation for the anisotropic components,
$\mathbf{\hat{D}}\pi_{n,l}=-\frac{\pi_{n,l}}{\tau_{R}}+c_{n,l}^{2}\frac{\pi_{1,1}}{\tau}+\left[c_{n,l}^{2}(\varepsilon+P)+\kappa_{n,l}-\frac{3}{G_{n+3,1}(z)}\left(G_{n+3,l}(z)-\frac{(2l-n)}{(2l+3)}G_{n+3,l+1}(z)\right)\rho_{n,1}^{\rm eq}\right]~\frac{1}{\tau}\,.$ (201)
Evolution of $\Pi$
The bulk pressure is related to the anisotropic moments $\Pi=-\frac{m^{2}}{3}\pi_{-1,0}$. The evolution equation for $\pi_{-1,0}$ is,
$\mathbf{\hat{D}}\pi_{-1,0}=-\frac{\pi_{-1,0}}{\tau_{R}}-\mathbf{\hat{D}}\rho^{\rm eq}_{-1,0}(\tau)$ (202)
$\mathbf{\hat{D}}\pi_{-1,0}=-\frac{\pi_{-1,0}}{\tau_{R}}+c_{-1,0}^{2}\frac{\pi_{1,1}}{\tau}+\left[c_{-1,0}^{2}(\varepsilon+P)+\kappa_{-1,0}-\frac{3}{G_{2,1}(z)}\left(G_{2,0}(z)-\frac{(2\cdot 0+1)}{(2\cdot 0+3)}G_{2,1}(z)\right)\rho_{-1,1}^{\rm eq}\right]~\frac{1}{\tau}\,.$ (203)
Multiplying both sides by $m^{2}/3$, we get the equation for $\Pi$,
$\mathbf{\hat{D}}\Pi$ $=-\frac{\Pi}{\tau_{R}}-\frac{m^{2}}{3}c_{-1,0}^{2}\frac{\pi_{1,1}}{\tau}-\left[\frac{m^{2}}{3}c_{-1,0}^{2}(\varepsilon+P)+\frac{m^{2}}{3}\kappa_{-1,0}-\frac{3}{G_{2,1}(z)}\left(G_{2,0}(z)-\frac{(2\cdot 0+1)}{(2\cdot 0+3)}G_{2,1}(z)\right)(\frac{m^{2}}{3}\rho_{-1,1}^{\rm eq})\right]~\frac{1}{\tau}$
$=-\frac{\Pi}{\tau_{R}}-\frac{m^{2}}{3}c_{-1,0}^{2}\frac{\pi_{1,1}}{\tau}-\left[\frac{m^{2}}{3}c_{-1,0}^{2}(\varepsilon+P)+\frac{m^{2}}{3}\kappa_{-1,0}-\left(\frac{G_{2,0}(z)}{G_{2,1}(z)}-\frac{1}{3}\right)\frac{G_{2,1}}{3G_{2,0}}(\varepsilon-3P)\right]~\frac{1}{\tau}$
$=-\frac{\Pi}{\tau_{R}}-\frac{m^{2}}{3}c_{-1,0}^{2}\frac{\pi_{1,1}}{\tau}-\left[\frac{m^{2}}{3}c_{-1,0}^{2}(\varepsilon+P)+\frac{m^{2}}{3}\kappa_{-1,0}-\frac{1}{3}\left(1-\frac{1}{3}\frac{G_{2,0}(z)}{G_{2,1}(z)}\right)(\varepsilon-3P)\right]~\frac{1}{\tau}$
We now relate $\frac{m^{2}}{3}c_{-1,0}^{2}$ to $c_{s}^{2}=c_{1,1}^{2}$ using the relations from Eq(189) and Eq.(116) . We have the explicit expression,
$\frac{m^{2}}{3}c_{-1,0}^{2}$ $=\frac{m^{2}}{3}~\frac{T^{-1-1}}{(2\cdot 0+1)}\frac{G_{-1+4,l}G_{3,0}-G_{-1+3,0}G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$
$=\frac{1}{3}\frac{m^{2}}{T^{2}}~\frac{G_{3,0}G_{3,0}-G_{2,0}G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$
$=\frac{1}{3}\frac{(z^{2}G_{3,0})G_{3,0}-(z^{2}G_{2,0})G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$ (204)
$=\frac{1}{3}\frac{(G_{5,0}-G_{5,1})G_{3,0}-(G_{4,0}-G_{4,1})G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$
$=\frac{1}{3}\frac{(G_{5,0}G_{3,0}-G_{4,0}^{2})-(G_{5,1}G_{3,0}-G_{4,1}G_{4,0})}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$
$=\frac{1}{3}\left[1-\frac{G_{5,1}G_{3,0}-G_{4,1}G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}\right]$ (205)
$=\frac{1}{3}-c_{s}^{2}\,.$ (206)
Similarly for $\frac{m^{2}}{3}\kappa_{-1,0}$ we have
$\frac{m^{2}}{3}\kappa_{-1,0}^{2}$ $=\frac{m^{2}}{3}~{\rm n}\frac{T^{-1}}{(2\cdot 0+1)}\frac{G_{5,0}G_{-1+3,0}-G_{4,0}G_{-1+4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$
$={\rm n}\frac{T}{3}\frac{G_{5,0}(G_{4,0}-G_{4,1})-G_{4,0}(G_{5,0}-G_{5,1})}{G_{5,0}G_{3,0}-G_{4,0}^{2}}$
$=-{\rm n}\frac{T}{3}\frac{G_{5,0}G_{4,1}-G_{5,1}G_{4,0}}{G_{5,0}G_{3,0}-G_{4,0}^{2}}=-\kappa$ (207)
From this we get
$\mathbf{\hat{D}}\Pi=-\frac{\Pi}{\tau_{R}}-\left[\frac{1}{3}-c_{s}^{2}\right]\frac{\pi_{1,1}}{\tau}-\left[\left(\frac{1}{3}-c_{s}^{2}\right)(\varepsilon+P)-\kappa-\frac{1}{3}\left(1-\frac{1}{3}\frac{G_{2,1}(z)}{G_{2,0}(z)}\right)(\varepsilon-3P)\right]\frac{1}{\tau}$ (208)
Using
$\mathbf{\hat{D}}\Pi=\partial_{\tau}\Pi+\frac{1}{\tau}\Pi+\frac{m^{2}}{3}\frac{1}{\tau}\pi_{-1,1}$ (209)
we get the final expression,
$\partial_{\tau}\Pi+\frac{\Pi}{\tau_{R}}=-\left[\frac{4}{3}-c_{s}^{2}\right]\frac{\Pi}{\tau}+\left[\frac{1}{3}-c_{s}^{2}\right]\frac{\pi}{\tau}-\left[\left(\frac{1}{3}-c_{s}^{2}\right)(\varepsilon+P)-\kappa-\frac{1}{3}\left(1-\frac{1}{3}\frac{G_{2,1}(z)}{G_{2,0}(z)}\right)(\varepsilon-3P)\right]\frac{1}{\tau}-\frac{m^{2}}{3}\frac{1}{\tau}\pi_{-1,1}$ (210)
Evolution of $\pi$
@@
$\mathbf{\hat{D}}\pi_{1,1}$ $=-\frac{\pi_{1,1}}{\tau_{R}}+c_{s}^{2}\frac{\pi_{1,1}}{\tau}+\left[c_{s}^{2}(\varepsilon+P)+\kappa-3\left(1-\frac{1}{5}~\frac{G_{4,2}}{G_{4,1}}\right)P\right]\frac{1}{\tau}\,.$ (211)
Now,